OpenStudy (anonymous):

Convert 0.49% to a decimal

OpenStudy (callisto):

1% = 1/100 = 0.01 0.49% = 0.49/100 = ...?

OpenStudy (anonymous):

0.49% = 0.49/100 = 0.0049

OpenStudy (unklerhaukus):

\[0.49\%=\frac{0.49}{100}=\frac{0.049}{10}={0.0049}\]